Write down the equation described in these word problems. Don't solve the problem. Just write down the equation.
#18 Eight times the sum of two and some number is 120
#19 The product of 10, and a number decreased by 5, is 200.
#20 Twelve subtracted by twice a number is four less than the same number

Answer by mananth(16949)

Eight times the sum of two and some number is 120
8(n+2)=120
The product of 10, and a number decreased by 5, is 200.
10n-5 =200
Twelve subtracted by twice a number is four less than the same number
2n-12 = n-4